OpenStudy (anonymous):

Correct please!!! A soccer team wins 65% of its matches, and 15% of its matches end in a draw. If the team is scheduled to play 20 matches, about how many matches is it expected to lose? I think it is 4...yes or no ???

OpenStudy (cp9454):

hell yeah

OpenStudy (anonymous):

can you show me how?

OpenStudy (owlcoffee):

Well, we will use some of understanding of the question at hand. Since the soccer team WINS 65% of it's matches and 15% of them end in DRAW, meaning that 20% ends up in defeat. The team is scheduled to play 20 matches, so that'll mean that we can discard 80% as NOT LOSING. therefore, 20% of the matches are possible defeats. That is the same as multiplying 20 by 1/5, since 20% is a fifth of 100%. (100/20=5): \[20 . \frac{ 1 }{ 5 }=\frac{ 20 }{ 5 }=4\]

OpenStudy (anonymous):

thank you for explaining. when i think I get something I realize i really have no idea.